# Self-supervised Representation Learning from Random Data Projectors

### v1

@@ Abstract
Self-supervised representation learning (SSRL) has advanced considerably by exploiting the transformation invariance assumption under artificially designed data augmentations. While augmentation-based SSRL algorithms push the boundaries of performance in computer vision and natural language processing, they are often not directly applicable to other data modalities, and can conflict with application-specific data augmentation constraints. This paper presents an SSRL approach that can be applied to any data modality and network architecture because it does not rely on augmentations or masking. Specifically, we show that high-quality data representations can be learned by reconstructing random data projections. We evaluate the proposed approach on a wide range of representation learning tasks that span diverse modalities and real-world applications. We show that it outperforms multiple state-of-the-art SSRL baselines. Due to its wide applicability and strong empirical results, we argue that learning from randomness is a fruitful research direction worthy of attention and further study. We provide code for our experiments at https://github.com/layer6ai-labs/lfr.
@@ 1 Introduction
While these examples show modality limitations, a more subtle problem is that even standard augmentations can conflict with application-specific constraints. For example, pathology images of stained tissue samples have low color variation (Shen et al., 2022; Kang et al., 2023), so naïve color-jittering produces unnatural augmentations (see Fig. 1). Such incorrect images may be inappropriate and unsafe for use in critical medical imaging applications (Elgendi et al., 2021). For
@@
time series with high periodicity, such as online monitoring data, random shifting augmentations can create identical augmented views that are not useful for pretext tasks (Eldele et al., 2021; Zhang & Ma, 2022). In tabular settings where relatively few options for augmentations are available, random noise addition or random swapping of features between training examples can easily produce unrealistic examples. As a concrete example from particle physics, consider energies and momenta of interacting particles collected experimentally (Brehmer & Cranmer, 2020). Since the incoming and outgoing energies are constrained by the laws of physics, only certain combinations are possible to observe. Tabular augmentations would produce unphysical combinations, harming the consistency of learned representations and leading to unpredictable failures on downstream tasks (Shorten & Khoshgoftaar, 2019).
@@ 3.1 Pretext Task: Multi-objective Learning from Randomness
We tackle the training instability issue by adopting the classic Expectation-Maximization (EM) method. Considering the MLE lower-bound (Objective (2)), optimizing the representation model $f_{\theta}$ is an E-step that repositions the posterior distribution of data $p(\mathcal{Z}|\mathcal{X})$ given fixed log-likelihood estimation modules $\log p\big{(}\mathbf{y}_{i}^{(k)}\mid\mathbf{z}_{i}\big{)}$. For the M-step, the representation distribution is fixed and we optimize the predictor heads $H_{\Phi}$. Hence, in this work we train the proposed SSRL model by alternating steps:
@@ 3.3 Diversity Encouragement on Random Data Projectors
By flattening the matrix $A^{(k)}$ and again normalizing, we obtain a vector $\mathbf{a}^{(k)}\in\mathbb{R}^{m^{2}\times 1}$, which acts as the signature of the $k$’th projector with respect to the batch.
Finally, to select $K$ target models from the $N$ candidates, we search for a subset that maximizes the following binary constraint optimization problem involving matrices $\tilde{A}$ made from $K$ stacked vectors $\mathbf{a}^{(k)}$,
@@ 4.1 Datasets
Computer vision: We tested on Kvasir (Pogorelov et al., 2017), a medical image dataset consisting of 8,000 images of the gastrointestinal tract. There are eight balanced classes including seven disease types as well as healthy images. We followed (Kalra et al., 2023) to resize all images to 100$\times$80 pixels and split the data into 6,000 images for training and 2,000 for testing.
@@ 4.2 Implementations
Evaluation: All the downstream tasks in our study are treated as classification problems. To evaluate the quality of the pre-trained representations, we employed supervised classifiers that are specific to each dataset. For the MIMIC-III dataset we utilized a MLP classifier (Yèche et al., 2021). For tabular datasets, we used logistic regression, similar to the approach in STab (Hajiramezanali et al., 2022). For the remaining datasets, a linear classifier was employed. The classifiers were trained on the frozen representations of the training set and evaluated on the test set, which is the most commonly used protocol. Accuracy is our primary evaluation metric, except for MIMIC-III where we adopted linearly weighted Cohen’s Kappa as in (Yèche et al., 2021), with higher values indicating better agreement. To ensure the robustness of our results, we conducted multiple random runs and report the mean and standard deviation, using 5 runs for tabular datasets and 3 runs for others.
Model architectures: Regarding the model architectures, we adopted similar backbone encoders as previous works. For the HAR and Epilepsy datasets, we utilized the same 3-block convolutional layers as TS-TCC (Wang et al., 2017). For the MIMIC-III dataset, we employed the Temporal Convolutional Network used by NCL (Yèche et al., 2021). For the Tabular datasets, we used 4-layer MLPs, following the approach in SCARF (Bahri et al., 2022). For Kvasir, we employed the ResNet18 architecture (He et al., 2016). To avoid domain-specific projector design, in each case the random projectors reuse the architecture from the encoder, but are scaled down. Complete details are in Appendix C.
@@ 4.3 Performance Across Data Modalities
This experimental result reflects our hypothesis – it is feasible to learn high-quality data representations across all modalities tested by predicting random data projections. LFR shows comparatively good performance on domains where semantic-preserving augmentations are difficult to create. In SSRL, fine-tuning evaluation is often also performed, so we conducted this analysis in Appendix D.3, along with information on computing resources.
@@ 5 Conclusion
This paper presents a novel self-supervised representation learning framework that is both modality-agnostic and application-agnostic. Our proposed framework utilizes random projectors to learn representations from unlabelled data, demonstrating excellent performance across various modalities and applications, particularly in situations where robust augmentation pipelines are not yet established.
The LFR technique is best suited to situations where the data cannot be reasonably augmented (due to the lack of domain knowledge). Although surprising, this situation occurs frequently in critical application domains, such as healthcare as we have highlighted. For general applications, however, if one knows the application domain well with adequate intuition around sensible data augmentations, using contrastive-learning-based SSRL is still likely to outperform random projectors. Overall, we treat LFR as a great complement to SSRL literature to fill the gap of data augmentation-free SSRL that satisfies the needs of many crucial applications.
@@ C.2 Neural Network Architectures
HAR/Epilepsy:
@@
Kvasir:
For LFR, we used ResNet18 (He et al., 2016) as the representation model, with an output dimension of 2048 for all datasets. The predictors are 4-layer ReLU networks with a hidden dimension of 256. For the random projector networks, we adopted a 4-layer CNN of channels $[3,8,16,32]$, each layer is followed by a ReLU, and every two layers are followed by max-pooling. A linear layer is used to map the output to 2048 dimensions. For the contrastive baselines, we adopted the same representation model for a fair comparison and used 3-layer ReLU networks with hidden dimension 256 as projectors. To create the contrastive views for Kvasir, we employed the augmentation set adopted by SimSiam (Chen et al., 2020) for ImageNet (RandomResizedCrop, RandomHorizontalFlip, ColorJitter, RandomGrayscale, GaussianBlur). We applied the same set of augmentations to the implementation of DIET (Balestriero, 2023). All other settings are the same as with the ResNet18 model.
@@
Table 4: Details on LFR architectural parameters
HAR/Epilepsy 6 Pytorch Default Three-block CNN Two-block CNN
Income/Theorem 6 Pytorch Default Four-layer MLP Two-layer MLP
HEPMASS 6 Pytorch Default Four-layer MLP Two-layer MLP
MIMIC-III 10 Pytorch Default Five-block TCN Three-block TCN
@@ C.3 Details of Training Settings
LFR training settings: Table 5 summarizes all the training settings used for LFR, while Table 6 outlines the evaluation settings used for downstream tasks. We used a Logistic Regression classifier for all tabular datasets including Income, Theorem, and HEPMASS, while for MIMIC-III we used a MLP network to predict the length of stay following (Yèche et al., 2021). For the remaining datasets, we followed prior works such as TS-TCC (Eldele et al., 2021), SimSiam (Chen et al., 2020), and BYOL (Le-Khac et al., 2020) by using a linear classifier for classification tasks.
@@
HAR/Epilepsy Adam 128 3e-4 $\beta$=(0.9, 0.999), wd=3e-4 Train epochs = 200, Predictor epochs = 5
@@
Baseline training settings: All self-supervised baselines adopt the same training setting as LFR unless stated otherwise. For DIET with MIMIC-III, we used batch size 512 and trained for 2000 steps with 10 warmup epochs. We reserved 5000 epochs for training the autoencoder on MIMIC-III with 500 warmup epochs. For other self-supervised methods with MIMIC-III, we also added 60 warmup epochs. We summarize the training settings of supervised baselines in Table 7.
@@
HAR/Epilepsy Adam 128 3e-4 $\beta$=(0.9, 0.999), wd=3e-4 500 None
@@ Appendix D Additional Experimental Results
D.1 Embedding Dimensions
@@
D.2 Sample size for projector selection
@@
D.3 Fine-tuning performance
@@
Through the fine-tuning process, all methods exhibit more comparable performance across the datasets. LFR still achieved the best performance on a majority of the datasets we used, although with overlapping error bars to other methods in those cases.
@@
D.4 Random Projectors Visualization
To illustrate the behaviour of individual random projections, and the role of diversity selection, we trained LFR on a small image-based dataset, CIFAR-10222CIFAR-10 was chosen for this qualitative study rather than the other datasets we used because it has human-interpretable features and visualizations., and examined the nearest-neighbour representations for different projectors. First we selected a query image at random, then for each of 100 randomly initialized projectors, we encoded all training images, and found the nearest encoded representations to the query. Distances in embedding space are measured by cosine similarity. We then visually compared the nearest neighbours that were selected by pairs of projectors which were deemed most diverse or similar according to the DPP criterion from Equation (9). The pairs with the highest diversity (Diverse 1 and Diverse 2) and the highest similarity (Similar 1 and Similar 2) are shown in Figure 8.
@@ Appendix E Computational Resources and Time Spent
The time series experiments with HAR and Epilepsy were conducted on a Tesla V100 GPU with 32 GB of memory, except for TS-TCC which was conducted on a TITAN V with 12 GB of memory. The experiments took a total of 102 GPU hours, including all baseline experiments. The MIMIC-III experiments were conducted with an NVIDIA A100 GPU with 40GB of memory, except for TS-TCC which was again conducted on a TITAN V with 12 GB of memory, and cost 608 GPU hours, including all baseline experiments. The Kvasir experiments were conducted using a Tesla V100 GPU with 32 GB of memory, and they took a total of 1095 GPU hours, including all baseline experiments. The tabular dataset experiments with Income, Theorem, and HEPMASS were conducted on an NVIDIA TITAN V GPU with 12 GB of memory. The experiments took a total of 70 GPU hours, including all baseline experiments.

### v2

@@ Abstract
Self-supervised representation learning (SSRL) has advanced considerably by exploiting the transformation invariance assumption under artificially designed data augmentations. While augmentation-based SSRL algorithms push the boundaries of performance in computer vision and natural language processing, they are often not directly applicable to other data modalities, and can conflict with application-specific data augmentation constraints. This paper presents an SSRL approach that can be applied to any data modality and network architecture because it does not rely on augmentations or masking. Specifically, we show that high-quality data representations can be learned by reconstructing random data projections. We evaluate the proposed approach on a wide range of representation learning tasks that span diverse modalities and real-world applications. We show that it outperforms multiple state-of-the-art SSRL baselines. Due to its wide applicability and strong empirical results, we argue that learning from randomness is a fruitful research direction worthy of attention and further study.
@@ 1 Introduction
While these examples show modality limitations, a more subtle problem is that even standard augmentations can conflict with application-specific constraints. For example, pathology images of stained tissue samples have low color variation (Shen et al., 2022; Kang et al., 2023), so naïve color-jittering produces unnatural augmentations (see Fig. 1). Such incorrect images may be inappropriate and unsafe for use in critical medical imaging applications (Elgendi et al., 2021). For time series with high periodicity, such as online monitoring data, random shifting augmentations can create
@@
identical augmented views that are not useful for pretext tasks (Eldele et al., 2021; Zhang & Ma, 2022). In tabular settings where relatively few options for augmentations are available, random noise addition or random swapping of features between training examples can easily produce unrealistic examples. As a concrete example from particle physics, consider energies and momenta of interacting particles collected experimentally (Brehmer & Cranmer, 2020). Since the incoming and outgoing energies are constrained by the laws of physics, only certain combinations are possible to observe. Tabular augmentations would produce unphysical combinations, harming the consistency of learned representations and leading to unpredictable failures on downstream tasks (Shorten & Khoshgoftaar, 2019).
@@ 3.1 Pretext Task: Multi-objective Learning from Randomness
We tackle the training instability issue by adopting the classic Expectation-Maximization (EM) method. Considering the MLE lower-bound (Objective (2)), optimizing the representation model $f_{\theta}$ is an E-step that repositions the posterior distribution of data $p(\mathcal{Z}|\mathcal{X})$ given fixed log-likelihood estimation modules $\log p\big{(}\mathbf{y}_{i}^{(k)}\mid\mathbf{z}_{i}\big{)}$. For the M-step, the representation distribution is fixed and we optimize the predictor heads $H_{\Phi}$. Details of the derivation can be found in Appendix G. Hence, in this work we train the proposed SSRL model by alternating steps:
@@ 3.3 Diversity Encouragement on Random Data Projectors
By flattening the matrix $A^{(k)}$ and again normalizing, we obtain a vector $\mathbf{a}^{(k)}\in\mathbb{R}^{m^{2}\times 1}$, which acts as the signature of the $k$’th projector with respect to the batch. Finally, to select $K$ target models from the $N$ candidates, we search for a subset that maximizes the following binary constraint optimization problem involving matrices $\tilde{A}$ made from $K$ stacked vectors $\mathbf{a}^{(k)}$,
@@ 4.1 Datasets
Computer vision: We tested on Kvasir (Pogorelov et al., 2017), a medical image dataset consisting of 8,000 images of the gastrointestinal tract. There are eight balanced classes including seven disease types as well as healthy images. We followed (Kalra et al., 2023) to resize all images to 100$\times$80 pixels and split the data into 6,000 images for training and 2,000 for testing. We provide further results on CIFAR10 in Appendix D.6.
@@ 4.2 Implementations
Evaluation: All the downstream tasks in our study are treated as classification problems. To evaluate the quality of the pre-trained representations, we employed supervised classifiers that are specific to each dataset. For the MIMIC-III dataset we utilized a MLP classifier (Yèche et al., 2021). For tabular datasets, we used logistic regression, similar to the approach in STab (Hajiramezanali et al., 2022). For the remaining datasets, a linear classifier was employed. The classifiers were trained on the frozen representations of the training set and evaluated on the test set, which is the most commonly used protocol. We also include finetuning results in Appendix D.4 and transfer learning results on Time Series in Appendix D.1. Accuracy is our primary evaluation metric, except for MIMIC-III where we adopted linearly weighted Cohen’s Kappa as in (Yèche et al., 2021), with higher values indicating better agreement. To ensure the robustness of our results, we conducted multiple random runs and report the mean and standard deviation, using 5 runs for tabular datasets and 3 runs for others.
Model architectures: Regarding the model architectures, we adopted similar backbone encoders as previous works. For the HAR and Epilepsy datasets, we utilized the same 3-block convolutional layers as TS-TCC (Wang et al., 2017). For the MIMIC-III dataset, we employed the Temporal Convolutional Network used by NCL (Yèche et al., 2021). For the Tabular datasets, we used 4-layer MLPs, following the approach in SCARF (Bahri et al., 2022). For Kvasir, we employed the ResNet18 architecture (He et al., 2016). To avoid domain-specific projector design, in each case the random projectors reuse the encoder architecture, but are scaled down. Complete details are in Appendix C.
@@ 4.3 Performance Across Data Modalities
This experimental result reflects our hypothesis – it is feasible to learn high-quality data representations across all modalities tested by predicting random data projections. LFR shows comparatively good performance on domains where semantic-preserving augmentations are difficult to create.
@@ 5 Conclusion
This paper presents a novel self-supervised representation learning framework that is both modality-agnostic and application-agnostic. Our proposed framework utilizes random projectors to learn representations from unlabelled data, demonstrating excellent performance across various modalities and applications, particularly in situations where robust augmentation pipelines are not yet established. The LFR technique is best suited to situations where the data cannot be reasonably augmented (due to the lack of domain knowledge). Although surprising, this situation occurs frequently in critical application domains, such as healthcare as we have highlighted. For general applications, however, if one knows the application domain well with adequate intuition around sensible data augmentations, using contrastive-learning-based SSRL is still likely to outperform random projectors. Overall, we treat LFR as a great complement to SSRL literature to fill the gap of data augmentation-free SSRL that satisfies the needs of many crucial applications.
@@ C.1 Dataset Summary
Fault Diagnosis (App. D.1) Time series Motor current signals 8184 2728 Multi-class classification (3) Accuracy
@@
CIFAR10 (App. D.6) Image Natural Images 50000 10000 Multi-class classification (10) Accuracy
@@ C.2 Neural Network Architectures
Table 4: Details on LFR architectural parameters
HAR/Epilepsy/Fault Diagnosis (App D.1) 6 Pytorch Default Three-block CNN Two-block CNN
Income/Theorem 6 Pytorch Default Four-layer MLP Two-layer MLP
HEPMASS 6 Pytorch Default Four-layer MLP Two-layer MLP
MIMIC-III 10 Pytorch Default Five-block TCN Three-block TCN
CIFAR10 (App D.6) 40 $\beta$ Initializtion + Weight Dropout ResNet18 Four-layer CNN
HAR/Epilepsy/Fault Diagnosis:
@@
Kvasir/CIFAR10:
For LFR, we used ResNet18 (He et al., 2016) as the representation model, with an output dimension of 2048 for all datasets. For CIFAR10 we followed SimSiam (Chen et al., 2020) to use the CIFAR variant of ResNet18. The predictors are 4-layer ReLU networks with a hidden dimension of 256. For the random projector networks, we adopted a 4-layer CNN of channels $[3,8,16,32]$, each layer is followed by a ReLU, and every two layers are followed by max-pooling. A linear layer is used to map the output to 2048 dimensions. For the contrastive baselines, we adopted the same representation model for a fair comparison and used 3-layer ReLU networks with hidden dimension 256 as projectors. To create the contrastive views, we employed the augmentation set adopted by SimSiam Chen et al. (2020) for CIFAR (RandomResizedCrop, RandomHorizontalFlip, ColorJitter, RandomGrayscale) and added GaussianBlur for Kvasir. We applied the same set of augmentations to the implementation of DIET (Balestriero, 2023). All other settings are the same as with the ResNet18 model.
@@ C.3 Details of Training Settings
LFR training settings: Table 5 summarizes all the training settings used for LFR, while Table 6 outlines the evaluation settings used for downstream tasks. We used a Logistic Regression classifier for all tabular datasets including Income, Theorem, and HEPMASS, while for MIMIC-III we used a MLP network to predict the length of stay following (Yèche et al., 2021). For the remaining datasets, we followed prior works such as TS-TCC (Eldele et al., 2021), SimSiam (Chen et al., 2020), and BYOL (Grill et al., 2020) by using a linear classifier for classification tasks.
@@
HAR/Epilepsy/Fault Diagnosis (App  D.1) Adam 128 3e-4 $\beta$=(0.9, 0.999), wd=3e-4 Train epochs = 200, Predictor epochs = 5
@@
CIFAR10 (App D.6) SGD 512 3e-2, cosine decay momentum=0.9, wd=5e-4 Train epochs = 400, Predictor epochs = 5
@@
CIFAR10 (App D.6) SGD 512 0.2, cosine decay momentum=0.9, wd=0 100
Baseline training settings: All self-supervised baselines adopt the same training setting as LFR unless stated otherwise. For DIET with CIFAR10, we followed the original paper to train 5000 epochs with 10 linear warmup epochs and then cosine decay. For DIET with MIMIC-III, we used batch size 512 and trained for 2000 steps with 10 warmup epochs. We reserved 5000 epochs for training the autoencoder on MIMIC-III with 500 warmup epochs. For other self-supervised methods with MIMIC-III, we also added 60 warmup epochs. We summarize the training settings of supervised baselines in Table 7.
@@
HAR/Epilepsy/Fault Diagnosis (App  D.1) Adam 128 3e-4 $\beta$=(0.9, 0.999), wd=3e-4 500 None
@@
CIFAR10 (App D.6) SGD 512 3e-2, cosine decay momentum=0.9, wd=5e-4 800 Same as SimCLR and SimSiam
@@ Appendix D Additional Experimental Results
D.1 Transfer Learning Experiments
We follow the settings in prior work (Eldele et al., 2021) and perform transfer learning experiments on a set of four real-world Fault Diagnosis datasets (Ragab et al., 2020). Each dataset was collected under a different working condition and can be considered as a separate domain with different characteristics. As shown in Table 8, LFR is able to learn features on the source task that transfer well to related tasks, outperforming the supervised approach (with the same model architecture) on 10 out of 12 dataset pairs, TS-TCC on 8 pairs, and is significantly better than all other approaches in terms of average performance. Hence, we see that LFR is able to transfer representations in a tabular setting, at least when there is some similarity in the domain of the datasets used.
Table 8: Transfer learning experiments
[…]
Supervised (Transf.) 34.28 44.94 34.57 52.93 63.67 99.82 52.93 84.02 83.54 53.15 99.56 62.43 63.83
Supervised (CNN) 42.96 46.33 46.99 43.55 70.53 94.94 48.50 78.34 74.34 52.71 99.05 70.20 64.04
Rand Init (CNN) 79.55 68.80 79.95 78.96 58.39 81.34 70.60 84.93 80.02 80.10 80.13 57.77 75.05
TS-TCC (Transf.) 43.15 51.50 42.74 47.89 70.38 99.30 38.89 98.31 99.38 51.91 99.96 70.31 67.83
LFR (CNN) 90.51 81.63 93.40 75.99 75.48 88.64 69.32 80.54 88.34 78.92 87.90 75.26 82.16
D.2 Embedding Dimensions
@@
D.3 Sample size for projector selection
@@
D.4 Fine-tuning performance
@@
As shown in Table 10, through the fine-tuning process all methods exhibit more comparable performance across the datasets. LFR still achieved the best performance on a majority of the datasets we used, although with overlapping error bars to other methods in those cases.
@@
D.5 Loss function ablations
In the main paper, we used the batch-wise Barlow Twins loss as a divergence measure, however this specific choice of loss function is not required for our method, and many other choices of contrastive loss could also work. As an ablation, we also ran LFR with the InfoNCE loss on the Kvasir dataset and observed a slight decrease compared to the Barlow Twins loss ($72.6\pm 0.4$ compared with $74.9\pm 0.6$).
D.6 Performance on CIFAR10
The primary focus of LFR is domain-agnostic representation learning, mainly in cases where it is not clear how to usefully augment the data. As a result, we focused on tabular, time-series, and medical imaging applications in the main paper. Here, we explore the effectiveness of LFR on natural images with CIFAR10 data. As shown in Table 11, as one might expect, SimSiam and SimCLR achieve the best performance out of our baselines thanks to their extensively engineered pipeline of image augmentations. Two domain-agnostic approaches that do not use image augmentations, DIET and our LFR, fall into a second tier, outpacing the third tier of DACL, Autoencoder, and Random Init. This result is in stark contrast to what we found on Kvasir, where all methods ended up in a tight range, with LFR at the top. This suggests that the heavy augmentations of SimSiam and SimCLR really are specialized for natural images, and are not suitable for other domains within the image modality.
Table 11: Performance on CIFAR10
Method Supervised Random Init Autoencoder DIET SimSiam SimCLR DACL LFR
Accuracy $94.4\pm 0.2$ $34.7\pm 0.3$ $37.4\pm 0.3$ $69.3\pm 1.2$ $89.2\pm 0.1$ $86.7\pm 0.5$ $42.4\pm 2.7$ $64.3\pm 0.1$
D.7 Random Projectors Visualization
To illustrate the behaviour of individual random projections, and the role of diversity selection, we trained LFR on a small image-based dataset, CIFAR10222CIFAR10 was chosen for this qualitative study rather than the other datasets we used because it has human-interpretable features and visualizations., and examined the nearest-neighbour representations for different projectors. First we selected a query image at random, then for each of 100 randomly initialized projectors, we encoded all training images, and found the nearest encoded representations to the query. Distances in embedding space are measured by cosine similarity. We then visually compared the nearest neighbours that were selected by pairs of projectors which were deemed most diverse or similar according to the DPP criterion from Equation (9). The pairs with the highest diversity (Diverse 1 and Diverse 2) and the highest similarity (Similar 1 and Similar 2) are shown in Figure 8.
@@ Appendix E Computational Resources and Time Spent
The time series experiments with HAR and Epilepsy were conducted on a Tesla V100 GPU with 32 GB of memory, except for TS-TCC which was conducted on a TITAN V with 12 GB of memory. The experiments took a total of 102 GPU hours, including all baseline experiments. The MIMIC-III experiments were conducted with an NVIDIA A100 GPU with 40GB of memory, except for TS-TCC which was again conducted on a TITAN V with 12 GB of memory, and cost 608 GPU hours, including all baseline experiments. The Kvasir experiments were conducted using a Tesla V100 GPU with 32 GB of memory, and they took a total of 1095 GPU hours, including all baseline experiments. The tabular dataset experiments with Income, Theorem, and HEPMASS were conducted on an NVIDIA TITAN V GPU with 12 GB of memory. The experiments took a total of 70 GPU hours, including all baseline experiments. The CIFAR experiments were conducted on a cluster with a single NVIDIA P100 GPU with 12 GB of memory per experiment, and they took a total of 3020 GPU hours, including all baseline experiments.
Appendix F Note on computational efficiency of LFR
Our experiments across diverse datasets consistently reveal that LFR’s training time is often comparable to, or even notably shorter than the contrastive SSRL methods. For example, on Kvasir training SimCLR took around 9 hours, while LFR completed in just 2 hours on the same machine. This assessment of training time considers both CPU and GPU usage.
The efficiency in LFR’s training time can be attributed to two primary factors. First, existing SSRL methods rely on extensive augmentations that are performed on CPU and can lead to data loading bottlenecks. LFR has no such issue. CPU-based image augmentations account for a significant amount of SimCLR’s 9 hour training time on Kvasir. For datasets with less computationally intensive augmentations, training times tend to align more closely - on the HAR dataset, LFR concluded training in approximately 3 hours while SCARF required around 3.3 hours.
Second, multiple passes through predictor heads in LFR is not overly burdensome. Unlike contrastive frameworks that use double augmentation and encoder passes, LFR requires fewer expensive operations: one pass through each random projector before training, and one pass through the encoder followed by one pass through each small predictor during training. Given that the encoder is the largest network by far, LFR’s training time often proves less than SimCLR which uses multiple encoder passes.
Compared to the cost of training an encoder, the diversity selection procedure using the Fast Determinantal Point Process (Chen et al. 2018) is an almost negligible contribution to the total training time. To recap, to select $K$ diverse projectors we initialize $N$ small neural networks and for each we run one forward pass on a batch of training data of size $m$. Then, DPP is applied over matrices of size $m^{2}\times K$ constructed from the network outputs. As a concrete example, on the Income dataset training LFR for 100 epochs took 849 s on our machine, while the DPP selection took 3 s.
Appendix G Mathematical derivation
Here we provide more details on the derivation of Equations 2 through 4.
Consider the MLE objective in Equation 2. As the variable $\mathbf{z}_{i}$’s distribution could be intractable, we represent it in a general form by the distribution $q(\mathbf{z}_{i})$ such that
$\sum_{i}\sum_{k}\log\int_{\mathbf{z}_{i}}p\left(\mathbf{y}_{i}^{(k)}\Bigm{|}\mathbf{z}_{i}\right)p\left(\mathbf{z}_{i}\mid\mathbf{x}_{i}\right)d\mathbf{z}_{i}=\sum_{i}\sum_{k}\log\int_{\mathbf{z}_{i}}q(\mathbf{z}_{i})\frac{p\left(\mathbf{y}_{i}^{(k)}\Bigm{|}\mathbf{z}_{i}\right)p\left(\mathbf{z}_{i}\mid\mathbf{x}_{i}\right)}{q(\mathbf{z}_{i})}d\mathbf{z}_{i}.$
Using Jensen’s inequality, the lower bound of the objective is therefore
[…]
Hence, to maximize the variational lower-bound of the above equation with respect to the proposed distribution $q(\mathbf{z}_{i})$, the optimal solution is simply to let equality hold (as in the classic EM algorithm),
[…]
which, in turn, requires the following equation to hold,
$\frac{p\left(\mathbf{y}_{i}^{(k)}\Bigm{|}\mathbf{z}_{i}\right)p\left(\mathbf{z}_{i}\mid\mathbf{x}_{i}\right)}{q(\mathbf{z}_{i})}=C,$
where $C$ is a constant. Thus, the optimal solution for $q(\mathbf{z}_{i})$ is
$q(\mathbf{z}_{i})=\frac{p\left(\mathbf{y}_{i}^{(k)}\Bigm{|}\mathbf{z}_{i}\right)}{p\left(\mathbf{y}_{i}^{(k)}\Bigm{|}\mathbf{x}_{i}\right)}p\left(\mathbf{z}_{i}\mid\mathbf{x}_{i}\right).$
As both $p\big{(}\mathbf{y}_{i}^{(k)}\bigm{|}\mathbf{z}_{i}\big{)}$ and $p\big{(}\mathbf{y}_{i}^{(k)}\bigm{|}\mathbf{x}_{i}\big{)}$ are delta distributions with probability $1$ conditioned on deterministic functions modelled by $g^{(k)}$ and $h^{(k)}_{\phi}$ respectively, the optimal solution of $q(\mathbf{z}_{i})$ is simply $p(\mathbf{z}_{i}\bigm{|}\mathbf{x}_{i})$ given the conditions are satisfied. In other words, we need $p(\mathbf{z}_{i}\bigm{|}\mathbf{x}_{i})$ that can let $\mathbf{y}_{i}^{(k)}=h^{(k)}_{\phi}(\mathbf{z}_{i})$ for all $k$. Thus, the optimization is essentially an EM algorithm where we optimize $\theta$ and $\phi$ alternatively to gradually increase the conditional likelihood of Equation 2.